f(x) = 3x - 7 and g(x) = -2x - 6 Find ( f o g) (4)
please show work

Answer by jim_thompson5910(35256)


The notation (f o g)(x) is really the same as f(g(x))

So saying (f o g)(4) means we want to compute f(g(4))

First let's plug x = 4 into g(x), which is the inner function.
g(x) = -2x-6
g(4) = -2(4)-6
g(4) = -8-6
g(4) = -14

So f(g(4)) gets updated to f(-14) after we replace the "g(4)" with "-14"

Now let's compute f(-14)
f(x) = 3x-7
f(-14) = 3(-14)-7
f(-14) = -42-7
f(-14) = -49

The final answer is therefore -49